# doped: Python toolkit for robust and repeatable charged defect supercell calculations

### 12 Mar 2024 v1

@@
Defects are a universal feature of crystalline solids, dictating the key properties and performance of many functional materials. Given their crucial importance yet inherent difficulty in measuring experimentally, computational methods (such as DFT and ML/classical force-fields) are widely used to predict defect behaviour at the atomic level and the resultant impact on macroscopic properties. Here we report doped, a Python package for the generation, pre-/post-processing and analysis of defect supercell calculations. doped has been built to implement the defect simulation workflow in an efficient, user-friendly yet powerful and fully-flexible manner, with the goal of providing a robust general-purpose platform for conducting reproducible calculations of solid-state defect properties.
@@
The materials science sub-field of computational defect modelling has seen considerable growth in recent years, driven by the crucial importance of these species in functional materials and the major advances in computational methodologies and resources facilitating their accurate simulation. Software which enables researchers to efficiently and accurately perform these calculations, while allowing for in-depth target analyses of the resultant data, is thus of significant value to the community. Indeed there are many critical stages in the computational workflow for defects, which when performed manually not only consume significant researcher time and effort but also leave room for human error – particularly for newcomers to the field. Moreover, there are growing efforts to perform high-throughput investigations of defects in solids (Xiong et al., 2023), necessitating robust, user-friendly and efficient software implementing this calculation workflow.
Given this importance of defect simulations and the complexity of the workflow, a number of software packages have been developed with the goal of managing pre- and post-processing of defect calculations, including work on the HADES/METADISE codes from the 1970s (Parker et al., 2004), to more recent work from Kumagai et al. (2021), Broberg et al. (2018), Shen & Varley (2024), Neilson & Murphy (2022), Arrigoni & Madsen (2021), Goyal et al. (2017), M. Huang et al. (2022), Péan et al. (2017) and Naik & Jain (2018). [^1] Some of these packages are no longer maintained, not compatible with high-throughput architectures, and/or are closed-source/commercial packages. While each of these codes have their strengths, they do not include the full suite of functionality provided by doped – some of which is discussed below – nor adopt the same focus on user-friendliness (along with sanity-checking warnings & error catching) and efficiency with full flexibility and wide-ranging functionality, targeting expert-level users and newcomers to the field alike.
@@
doped is a Python software for the generation, pre-/post-processing and analysis of defect supercell calculations, as depicted in Figure 1. The design philosophy of doped has been to implement the defect simulation workflow in an efficient, reproducible, user-friendly yet powerful and fully-customisable manner, combining reasonable defaults with full user control for each parameter in the workflow. As depicted in Figure 1, the core functionality of doped is the generation of defect supercells and competing phases, writing calculation input files, parsing calculation outputs and analysing/plotting defect-related properties. This functionality and recommended usage of doped is demonstrated in the tutorials on the documentation website.
Figure 2: a. Average minimum periodic image distance, normalised by the ideal image distance at that volume (i.e. for a perfect close-packed face-centred cubic (FCC) cell), versus the number of primitive unit cells, for the supercell generation algorithms of doped, ASE and pymatgen. “SC” = simple cubic and “HCP” = hexagonal close-packed. b. Average performance of charge state guessing routines from pyCDT, pymatgen-analysis-defects, defectivator and doped, in terms of false positives and false negatives. Asterisk indicates that the pyCDT false negatives are underestimated as the majority of this test set used the guessed charge state ranges from pyCDT. “Ox.” = oxidation & “prob.” = probabilities. Example (c) Kumagai-Oba (extended Freysoldt-Neugebauer-Van-de-Walle; “eFNV”) finite-size correction plot, (d) defect formation energy diagram, (e) chemical potential / stability region, (f) Fermi level vs. annealing temperature, (g) defect/carrier concentrations vs. annealing temperature and (h) Fermi level / carrier concentration heatmap plots from doped. Data and code to reproduce these plots is provided in the docs/JOSS subfolder of the doped GitHub repository.
@@
  • Charge-state guessing: Defects in solids can adopt various electronic charge states. However, the set of stable charge states for a given defect is typically not known a priori and so one must choose a set of possible defect charge states to calculate – usually relying on some form of chemical intuition. In this regard, extremal defect charge states that are calculated but do not end up being stable can be considered ‘false positives’ or ‘wasted’ calculations, [^2] Note that unstable defect charge states which are intermediate between stable charge states (e.g. $X^{0}$ for a defect $X$ with a (+1/-1) negative-U level) should still be calculated and are not considered false positives. while charge states which are stable but were not calculated can be considered ‘false negatives’ or ‘missed’ calculations. doped builds on other routines which use known elemental oxidation states to additionally account for oxidation state probabilities, the electronic state of the host crystal and charge state magnitudes. Implementing these features in a simple cost function, we find a significant improvement in terms of both efficiency (reduced false positives) and completeness (reduced false negatives) for this charge state guessing step, as shown in Figure 2b. [^3] Given sufficient data, a machine learning model could likely further improve the performance of this charge state guessing step.
@@
  • Automated symmetry & degeneracy handling: doped automatically determines the point symmetry of both initial (un-relaxed) and final (relaxed) defect configurations, and computes the corresponding orientational (and spin) degeneracy factors. This is a key pre-factor in the defect concentration equation:
@@
    where $g$ is the product of all degeneracy factors, $N_{s}$ is the concentration of lattice sites for that defect, $E_{f}$ is the defect formation energy and $N_{D}$ is the defect concentration. $g$ can affect predicted defect/carrier concentrations by up to 2/3 orders of magnitude (Kavanagh, Scanlon, et al., 2022; Mosquera-Lois, Kavanagh, Klarbring, et al., 2023), and is often overlooked in defect calculations, partly due to the (previous) requirement of significant manual effort and knowledge of group theory.
@@
  • Thermodynamic analysis: doped provides a suite of flexible tools for the analysis of defect thermodynamics, including formation energy diagrams (Figure 2d), equilibrium & non-equilibrium Fermi level solving (Figure 2f), doping analysis (Figure 2g,h), Brouwer-type diagrams etc. These include physically-motivated (but tunable) grouping of defect sites, full inclusion of metastable states, support for complex system constraints, optimisation over high-dimensional chemical & temperature space and highly-customisable plotting.
  • Finite-size corrections: Both the isotropic Freysoldt (FNV) (Freysoldt et al., 2009) and anisotropic Kumagai (eFNV) (Kumagai & Oba, 2014) image charge corrections are implemented automatically in doped, with tunable sampling radii / sites (which may be desirable for e.g. layered materials), automated correction plotting (to visualise/analyse convergence; Figure 2c) and automatic sampling error estimation.
  • Reproducibility & tabulation: doped has been built to support and encourage reproducibility, with all input parameters and calculation results saved to lightweight JSON files. This allows for easy sharing of calculation inputs/outputs and reproducible analysis. Several tabulation functions are also provided to facilitate the quick summarising of key quantities as exemplified in the tutorials (including defect formation energy contributions, charge transition levels (with/without metastable states), symmetry, degeneracy and multiplicity factors, defect/carrier concentrations, chemical potential limits, dopability limits, doping windows…) to aid transparency, reproducibility, comparisons with other works and general analysis. The use of these tabulated outputs in supporting information of publications is encouraged.

### 11 Apr 2024 v2

@@
Defects are a universal feature of crystalline solids, dictating the key properties and performance of many functional materials. Given their crucial importance yet inherent difficulty in measuring experimentally, computational methods (such as DFT and ML/classical force-fields) are widely used to predict defect behaviour at the atomic level and the resultant impact on macroscopic properties. Here we report doped, a Python package for the generation, pre-/post-processing, and analysis of defect supercell calculations. doped has been built to implement the defect simulation workflow in an efficient and user-friendly – yet powerful and fully-flexible – manner, with the goal of providing a robust general-purpose platform for conducting reproducible calculations of solid-state defect properties.
@@
The materials science sub-field of computational defect modelling has seen considerable growth in recent years, driven by the crucial importance of these species in functional materials and the major advances in computational methodologies and resources facilitating their accurate simulation. Software which enables researchers to efficiently and accurately perform these calculations, while allowing for in-depth target analyses of the resultant data, is thus of significant value to the community. Indeed there are many critical stages in the computational workflow for defects, which when performed manually not only consume significant researcher time and effort but also leave room for human error – particularly for newcomers to the field. Moreover, there are growing efforts to perform high-throughput investigations of defects in solids (Broberg et al., 2023; Xiong et al., 2023; Yuan et al., 2024), necessitating robust, user-friendly, and efficient software implementing this calculation workflow.
Given this importance of defect simulations and the complexity of the workflow, a number of software packages have been developed with the goal of managing pre- and post-processing of defect calculations, including work on the HADES/METADISE codes from the 1970s (Parker et al., 2004), to more recent work from Kumagai et al. (2021), Broberg et al. (2018), Shen & Varley (2024), Neilson & Murphy (2022), Arrigoni & Madsen (2021), Goyal et al. (2017), M. Huang et al. (2022), Péan et al. (2017) and Naik & Jain (2018). [^1] Some of these packages are no longer maintained, not compatible with high-throughput architectures, and/or are closed-source/commercial packages. While each of these codes have their strengths, they do not include the full suite of functionality provided by doped – some of which is discussed below – nor adopt the same focus on user-friendliness (along with sanity-checking warnings and error catching) and efficiency with full flexibility and wide-ranging functionality, targeting expert-level users and newcomers to the field alike.
@@
doped is a Python package for the generation, pre-/post-processing, and analysis of defect supercell calculations, as depicted in Figure 1. The design philosophy of doped has been to implement the defect simulation workflow in an efficient, reproducible, and user-friendly – yet powerful and fully-customisable – manner, combining reasonable defaults with full user control for each parameter in the workflow. As depicted in Figure 1, the core functionality of doped is the generation of defect supercells and competing phases, writing calculation input files, parsing calculation outputs, and analysing/plotting defect-related properties. This functionality and recommended usage of doped is demonstrated in the tutorials on the documentation website.
[…]
@@
  • Charge-state Estimation: Defects in solids can adopt various electronic charge states. However, the set of stable charge states for a given defect is typically not known a priori, so one must choose a set of possible defect charge states to calculate – usually relying on some form of chemical intuition. In this regard, extremal defect charge states that are calculated but do not end up being stable can be considered ‘false positives’ or ‘wasted’ calculations, [^2] Note that unstable defect charge states which are intermediate between stable charge states (e.g. $X^{0}$ for a defect $X$ with a (+1/-1) negative-U level) should still be calculated and are not considered false positives. while charge states which are stable but were not calculated can be considered ‘false negatives’ or ‘missed’ calculations. doped builds on other routines which use known elemental oxidation states to additionally account for oxidation state probabilities, the electronic state of the host crystal and charge state magnitudes. Implementing these features in a simple cost function, we find a significant improvement in terms of both efficiency (reduced false positives) and completeness (reduced false negatives) for this charge state estimation, as shown in Figure 2b. [^3] Given sufficient data, a machine learning model could likely further improve the performance of this charge state estimation.
@@
  • Automated Symmetry & Degeneracy Handling: doped automatically determines the point symmetry of both initial (un-relaxed) and final (relaxed) defect configurations, and computes the corresponding orientational (and spin) degeneracy factors. This functionality is also offered in the form of standalone functions which do not require the defect calculations to have been generated/parsed with doped. This is a key pre-factor in the defect concentration equation:
@@
    where $g$ is the product of all degeneracy factors, $N_{s}$ is the concentration of lattice sites for that defect, $E_{f}$ is the defect formation energy and $N_{D}$ is the defect concentration. $g$ can affect predicted defect/carrier concentrations by up to two or three orders of magnitude (Kavanagh, Scanlon, et al., 2022; Mosquera-Lois, Kavanagh, Klarbring, et al., 2023), and is often overlooked in defect calculations, partly due to the (previous) requirement of significant manual effort and knowledge of group theory.
@@
  • Thermodynamic Analysis: doped provides a suite of flexible tools for the analysis of defect thermodynamics, including formation energy diagrams (Figure 2d), equilibrium & non-equilibrium Fermi level solving (Figure 2f), doping analysis (Figure 2g,h), Brouwer-type diagrams etc. These include physically-motivated (but tunable) grouping of defect sites, full inclusion of metastable states, support for complex system constraints, optimisation over high-dimensional chemical & temperature space and highly customisable plotting. In-depth examples are provided in the tutorials.
  • Finite-Size Corrections: Both the isotropic Freysoldt (FNV) (Freysoldt et al., 2009) and anisotropic Kumagai (eFNV) (Kumagai & Oba, 2014) image charge corrections are implemented automatically in doped, with tunable sampling radii / sites (which may be desirable for e.g. layered materials), automated correction plotting (to visualise/analyse convergence; Figure 2c), and automatic sampling error estimation.
  • Reproducibility & Tabulation: doped has been built to support and encourage reproducibility, with all input parameters and calculation results saved to lightweight JSON files. This allows for easy sharing of calculation inputs/outputs and reproducible analysis. Several tabulation functions are also provided to facilitate the quick summarising of key quantities as exemplified in the tutorials (including defect formation energy contributions, charge transition levels (with/without metastable states), symmetry, degeneracy and multiplicity factors, defect/carrier concentrations, chemical potential limits, dopability limits, doping windows…) to aid transparency, reproducibility, comparisons with other works, and general analysis. The use of these tabulated outputs in supporting information of publications is encouraged.